OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12)

OpenStudy (raffle_snaffle):

where would you start?

OpenStudy (ahsome):

Here is a hint: \[m=\frac{y_2-y_1}{x_2-x_1}\] Where \(m\) is the slope

OpenStudy (anonymous):

Y-Yo=m(X-Xo) Plug in the values for X and Y and solve for the slope. After you find the slope plug it into the slope intercept form y=mx+b where m is your slope, b is some value you begin at. Choose any x value you're given.

OpenStudy (anonymous):

(-2, 7) and (3, 12) m=1 y=x-2

OpenStudy (anonymous):

What is the y=mx + b format on the answer

OpenStudy (anonymous):

hELP SOMEONE